### v1

@@ Abstract
Hyperparameter optimization (HPO) plays a central role in the performance of deep learning models, yet remains computationally expensive and difficult to interpret, particularly for time-series forecasting. While Bayesian Optimization (BO) is a standard approach, it typically treats tuning tasks independently and provides limited insight into its decisions. Recent advances in large language models (LLMs) offer new opportunities to incorporate structured prior knowledge and reasoning into optimization pipelines. We introduce LLM-AutoOpt, a hybrid HPO framework that combines BO with LLM-based contextual reasoning. The framework encodes dataset meta-features, model descriptions, historical optimization outcomes, and target objectives as structured meta-knowledge within LLM prompts, using BO to initialize the search and mitigate cold-start effects. This design enables context-aware and stable hyperparameter refinement while exposing the reasoning behind optimization decisions. Experiments on a multivariate time series forecasting benchmark demonstrate that LLM-AutoOpt achieves improved predictive performance and more interpretable optimization behavior compared to BO and LLM baselines without meta-knowledge.
@@ 1 Introduction
AutoML is the process of automating the tasks involved in building and deploying machine learning models [27]. One of the most challenging components of AutoML is hyperparameter optimization (HPO), which is essential for developing robust and effective deep learning models. The goal of HPO is to identify optimal configurations that enhance model performance, stability, and interpretability. Traditional HPO methods such as grid search, random search, Bayesian optimization, and gradient-based approaches are commonly used to optimize hyperparameters and have demonstrated improvements in model accuracy [28]. However, these methods still struggle with large-scale data and complex models including deep neural networks, and are often computationally expensive [1, 2, 6]. AutoML enhances the hyperparameter optimization task by incorporating different learning paradigms such as meta-learning and reinforcement learning.
@@
Meta-learning enables hyperparameter optimization (HPO) methods to rapidly adapt to new tasks by leveraging experience from previously observed tasks, where each task is represented through informative meta-features [3]. By transferring prior optimization knowledge, meta-learning–based HPO significantly improves both sample efficiency and adaptability compared to conventional approaches that treat each task independently. Crucially, these methods mitigate the need to initialize the optimization process from scratch, instead warm-starting the search using insights derived from historical tasks [4, 8, 9, 10]. This paradigm has been successfully operationalized in several AutoML systems, most notably Auto-sklearn [5], demonstrating its practical effectiveness in real-world settings.
@@
As large language models (LLMs) have become increasingly integrated into machine learning research, they have also been applied to hyperparameter optimization (HPO). Prior work has leveraged LLMs to automate tasks such as neural architecture search and end-to-end ML pipeline construction [11, 12]. Owing to their strong in-context learning capabilities, LLMs can rapidly adapt to new tasks without explicit fine-tuning [13, 14]. This property has recently been extended to optimization settings, where LLMs are explored as alternatives or complements to traditional HPO methods. For example, Bayesian optimization has been reformulated as a natural language reasoning task, enabling LLMs to exploit their contextual reasoning abilities to recommend improved hyperparameter configurations [17]. More recent approaches, such as MetaLLMix [23], integrate meta-learning with LLM reasoning and explainable AI to transfer knowledge across tasks using historical experiment meta-datasets. In this framework, LLMs reason over structured meta-information to generate task-specific model and hyperparameter recommendations, with SHAP-based explanations enhancing transparency. While such methods have shown promising results for classification and regression tasks [22], their applicability to time-series forecasting remains largely unexplored due to the presence of strong temporal dependencies and evolving data distributions. Prior work has operationalized meta-learning for hyperparameter optimization through several complementary strategies. Feurer et al. [9] demonstrate that Bayesian optimization can be effectively warm-started by conditioning surrogate models on dataset meta-features, thereby guiding the search toward promising regions of the hyperparameter space. In parallel, In [8], the authors propose transferring high-performing configurations from meta-datasets of previous tuning runs to initialize optimization on new tasks. Extending these ideas, Lindauer and Hutter [10] show that incorporating historical benchmark data into warm-started, model-based configuration frameworks improves scalability while preserving optimization efficiency. Collectively, these methods highlight how meta-knowledge can be systematically exploited to accelerate convergence and improve solution quality relative to cold-start optimization.
@@
To address this, we propose LLM-AutoMOpt, a framework for automating hyperparameter optimization in deep learning models for time series forecasting. The framework follows an iterative optimization process in which Bayesian optimization is first employed to initialize the search space and provide informative prior configurations, which are subsequently refined through structured LLM-guided reasoning. The LLM operates on carefully constructed prompts enriched with meta-knowledge, including dataset characteristics, model descriptions, target loss objectives, and historical optimization outcomes. This structured meta-information enables the model to generate context-aware and performance-driven hyperparameter recommendations. A key challenge lies in prompt design, which must balance expressive reasoning with well-defined constraints to mitigate hallucination and ensure stable optimization behavior. Beyond identifying optimal configurations, LLM-AutoMOpt emphasizes transparency and interpretability, explicitly exposing the meta-knowledge and reasoning processes underlying its optimization decisions.
@@ 2 Related works
Hyperparameter optimization (HPO) seeks to automatically identify hyperparameter configurations that maximize a model’s performance on a given task, often at the cost of expensive and time-consuming model evaluations [27]. Early approaches such as grid search systematically enumerate predefined configurations, guaranteeing coverage of the search space but becoming computationally prohibitive as the number or range of hyperparameters grows [36]. Random search addresses this limitation by sampling configurations independently, frequently achieving superior efficiency—particularly when only a small subset of hyperparameters has a dominant effect on performance [37]. Bayesian optimization further advances HPO by learning a probabilistic surrogate of the objective function and selecting new configurations via an explicit exploration–exploitation trade-off, enabling faster convergence with significantly fewer evaluations [29]. Despite their success, these methods generally treat each optimization problem independently, failing to exploit knowledge from prior tasks [1]. This limitation reduces scalability and sample efficiency, especially in computationally demanding settings such as time-series forecasting, where training costs are high and model behavior is strongly data-dependent [2].
@@
To overcome this limitation, meta-learning has emerged as an effective paradigm for improving scalability and resource efficiency across a wide range of learning problems [30]. By leveraging knowledge acquired from previously solved tasks, meta-learning enables hyperparameter optimization to adapt more rapidly to new tasks, effectively avoiding repeated optimization from scratch. In particular, Bayesian optimization has been shown to benefit significantly from meta-learning–driven warm starts, where prior task information guides the initial search process [31, 32]. This combination has been successfully applied to complex attention-based architectures, leading to improved forecasting accuracy in energy systems compared to standalone optimization approaches [33]. Building on this idea, meta-Bayesian optimization learns informative surrogate model priors from related tasks, further accelerating convergence and improving optimization efficiency on new problems [38].
@@
Beyond Bayesian and meta-learning–based approaches, reinforcement learning (RL) has also been explored for hyperparameter optimization by formulating HPO as a sequential decision-making problem. In this setting, an RL agent iteratively selects hyperparameter configurations based on feedback from previous evaluations, learning a policy that aims to maximize expected performance while reducing the number of costly evaluations [34]. Building on this idea, meta-learning has been combined with RL to further improve efficiency. For example, MultiTaskPB2 leverages metadata from prior RL training runs to warm-start hyperparameter optimization and bias surrogate models for new tasks, achieving significantly faster and more effective optimization than standard methods [35].
@@
More recently, large language models (LLMs) have introduced a new paradigm for hyperparameter optimization by leveraging natural language understanding, contextual reasoning, and few-shot generalization capabilities [15, 16, 17, 18, 19]. In AgentHPO [15], an LLM-based agent autonomously analyzes task descriptions and historical trials, proposes new hyperparameter configurations, executes experiments, and iteratively refines its recommendations, forming a human-like and interpretable optimization loop that reduces manual intervention. LLMs have also been integrated with traditional HPO strategies to improve efficiency. For instance, coarse grid search can be guided by an LLM to identify and refine promising regions of the search space, effectively pruning unlikely configurations and reducing computational cost while maintaining or improving tuning quality [18]. Furthermore, LLMs have been combined with Bayesian optimization in frameworks such as LLAMBO, where the BO process is reformulated as a natural language reasoning task. This enables LLMs to propose and evaluate candidate configurations conditioned on past trials, supporting zero-shot warm-starting, improved surrogate modeling, and more focused exploration of promising regions, ultimately leading to faster convergence and superior hyperparameter tuning compared to conventional BO [17]. Beyond directly steering the optimization process, recent research has investigated how LLMs can be used as meta-learners that reason over collections of prior experiments. Frameworks such as MetaLLMix [22] construct structured meta-databases from historical training runs and rely on LLMs to synthesize transferable optimization knowledge across tasks. Related in-context meta-learning approaches [23] encode datasets and experimental settings as structured, human-readable metadata, enabling LLMs to infer suitable hyperparameter configurations in zero-shot or few-shot regimes.
@@
Unlike agent-based or BO-integrated LLM methods, these approaches emphasize knowledge transfer at the task level rather than iterative trial-by-trial optimization. However, they have primarily been evaluated on classification and regression benchmarks and do not explicitly address the domain-specific challenges of time-series forecasting, where strong temporal dependencies, heterogeneous sampling rates, seasonal dynamics, and high training costs require specialized meta-features and optimization strategies. To address these limitations, we propose LLM-AutoMOpt, a forecasting-specific HPO framework that tightly integrates Bayesian optimization for structured initialization with LLM-based in-context reasoning over rich temporal meta-knowledge. In contrast to prior work, LLM-AutoMOpt explicitly accounts for temporal structure and model dynamics, enabling more sample-efficient, interpretable, and adaptive hyperparameter optimization for deep sequence models such as LSTM, Bi-LSTM, and Transformer architectures.
3 Methodology
LLM-AutoOpt is a hyperparameter optimization framework that exploits the reasoning capabilities of large language models (LLMs) to recommend high-quality hyperparameter configurations for deep learning models in time series forecasting. In addition to improving optimization efficiency, the framework emphasizes transparency by making the optimization process interpretable. LLM-AutoOpt operates iteratively, progressively refining the LLM’s contextual understanding through the integration of rich meta-knowledge derived from multiple stages of the model development pipeline. This meta-knowledge is embedded within structured prompts, enabling the LLM to reason effectively over dataset characteristics, model properties, and optimization objectives. By combining informed contextual reasoning with iterative feedback, the framework not only identifies optimal hyperparameter settings but also exposes the underlying rationale guiding the optimization decisions. This section describes the construction of the meta-knowledge, the prompt design, and the overall optimization workflow.
@@ 3.1 Forecasting-Specific Meta-Knowledge
Time-Series Meta-Features.
@@
Meta-Feature Statistical Summarization.
Beyond providing raw meta-features, we derive higher-level statistical summaries that aggregate these descriptors into interpretable indicators of the dataset behavior. Examples include summarizing autocorrelation coefficients across multiple lags to characterize overall temporal dependence, aggregating trend and stationarity statistics (e.g., ADF statistics and trend slopes) to describe global non-stationarity, and combining volatility- and noise-related measures (e.g., coefficient of variation, residual variance, and outlier ratio) to estimate the effective noise level of the series. These summarized representations are injected into the prompt in human-readable form, enabling the LLM to reason over dataset characteristics at a semantic level level rather than directly interpreting raw numerical statistics. This Abstraction facilitates better generalization in decision-making and significantly reduces hallucinated reasoning by constraining the LLM’s inference process to validate dataset-derived indicators.
Bayesian Optimization Initialization.
Bayesian Optimization (BO) is employed as an initial exploration mechanism to efficiently probe the hyperparameter space. The best-performing configuration—defined as the trial achieving the minimum observed validation loss. It is treated as task-specific meta-knowledge. Conditioning the LLM on this configuration provides a strong inductive bias, enabling it to refine an already high-quality solution rather than reasoning from an uninformed prior. This strategy improves sample efficiency, reduces convergence time, and mitigates the risk of stagnation in suboptimal regions of the search space.
Model Specification.

### v2

@@ Abstract
Hyperparameter optimization (HPO) plays a central role in the performance of deep learning models, yet remains computationally expensive and difficult to interpret, particularly for time-series forecasting. Although Bayesian Optimization (BO) is a standard approach, it typically treats tuning tasks independently and provides limited insight into its decisions. Recent advances in large language models (LLMs) offer new opportunities for incorporating structured prior knowledge and reasoning into optimization pipelines. We introduce LLM-AutoOpt, a hybrid HPO framework that combines BO with LLM-based contextual reasoning to address these issues. The framework encodes dataset meta-features, model descriptions, historical optimization outcomes, and target objectives as structured meta-knowledge within LLM prompts, using BO to initialize the search and mitigate the cold-start effects. This design enables context-aware and stable hyperparameter refinement while exposing the reasoning behind the optimization decisions. Experiments on a multivariate time series forecasting benchmark demonstrate that LLM-AutoOpt achieves improved predictive performance and more interpretable optimization behavior compared to BO and LLM baselines without meta-knowledge.
keywords:
AutoML, Hyper-Parameter Optimization (HPO), LLMs, Meta-Learning, Context-Aware Optimization, Time series forecasting, Interpretability
@@ 1 Introduction
Auto Ml is the process of automating the tasks involved in building and deploying machine learning models [ibm_automl_2025]. Hyperparameter optimization (HPO) is one of the most challenging components of AutoML, and is essential for developing robust and effective deep learning models. The goal of HPO is to identify optimal configurations that enhance the model performance, stability, and interpretability. Traditional HPO methods, such as grid search, random search, Bayesian optimization, and gradient-based approaches, are commonly used to optimize hyperparameters and have demonstrated improvements in model accuracy  [bib1]. However, these methods still struggle with large-scale data and complex models, including deep neural networks, and are often computationally expensive  [bib2, bib3, bib4].AutoML enhances the hyperparameter optimization task by incorporating different learning paradigms such as meta-learning and reinforcement learning.
@@
Meta-learning enables hyperparameter optimization (HPO) methods to rapidly adapt to new tasks by leveraging experience from previously observed tasks, where each task is represented through informative meta-features  [bib5]. By transferring prior optimization knowledge, meta-learning–based HPO significantly improves both sample efficiency and adaptability compared to conventional approaches that treat each task independently. Crucially, these methods mitigate the need to initialize the optimization process from scratch, instead warm-starting the search using insights derived from historical tasks  [bib6, bib7, bib8, bib9].This paradigm has been successfully operationalized in several AutoML systems, most notablyAuto-sklearn  [bib11], demonstrates its practical effectiveness in real-world settings.
@@
As large language models (LLMs) have become increasingly integrated into machine learning research, they have also been applied to hyperparameter optimization (HPO). Prior work has leveraged LLMs to automate tasks such as neural architecture search and end-to-end ML pipeline construction  [bib12, bib13].Owing to their strong in-context learning capabilities, LLMs can rapidly adapt to new tasks without explicit fine-tuning [bib14, bib15].This property has recently been extended to optimization settings, where LLMs are explored as alternatives or complements to traditional HPO methods. For example, Bayesian optimization has been reformulated as a natural language reasoning task, enabling LLMs to exploit their contextual reasoning abilities to recommend improved hyperparameter configurations [bib16].More recent approaches, such as MetaLLMix  [bib17], integrate meta-learning with LLM reasoning and explainable AI to transfer knowledge across tasks using historical experiment meta-datasets. In this framework, LLMs reason over structured meta-information to generate task-specific model and hyperparameter recommendations, with SHAP-based explanations enhancing transparency. While such methods have shown promising results for classification and regression tasks [bib18], their applicability to time-series forecasting remains largely unexplored due to the presence of strong temporal dependencies and evolving data distributions. Prior work has operationalized meta-learning for hyperparameter optimization through several complementary strategies. Feurer et al. [bib9] demonstrate that Bayesian optimization can be effectively warm-started by conditioning surrogate models on dataset meta-features, thereby guiding the search toward promising regions of the hyperparameter space. In parallel, In  [bib8], the authors propose transferring high-performing configurations from metadatasets of previous tuning runs to initialize optimization on new tasks. Extending these ideas, Lindauer and Hutter  [bib10] show that incorporating historical benchmark data into warm-started, model-based configuration frameworks improves scalability while preserving optimization efficiency.Collectively, these methods highlight how meta-knowledge can be systematically exploited to accelerate convergence and improve solution quality relative to cold-start optimization.
@@
To address this, we propose LLM-AutoMOpt, a framework for automating hyperparameter optimization in deep learning models for time series forecasting.The framework follows an iterative optimization process in which Bayesian optimization is first employed to initialize the search space and provide informative prior configurations, which are subsequently refined through structured LLMguided reasoning. The LLM operates on carefully constructed prompts enriched with meta-knowledge, including dataset characteristics, model descriptions, target loss objectives, and historical optimization outcomes. This structured metainformation enables the model to generate context-aware and performance-driven hyperparameter recommendations. A key challenge lies in prompt design, which must balance expressive reasoning with well-defined constraints to mitigate hallucination and ensure stable optimization behavior. Beyond identifying optimal configurations, LLM-AutoMOpt emphasizes transparency and interpretability, explicitly exposing the meta-knowledge and reasoning processes underlying its optimization decisions.
@@ 2 Related Works
Hyperparameter optimization (HPO) seeks to automatically identify hyperparameter configurations that maximize a model’s performance on a given task, often at the cost of expensive and time-consuming model evaluations [ibm_automl_2025].Early approaches such as grid search systematically enumerate predefined configurations, guaranteeing coverage of the search space but becoming computationally prohibitive as the number or range of hyperparameters grows [bib19].Random search addresses this limitation by sampling configurations independently, frequently achieving superior efficiency, particularly when only a small subset of hyperparameters has a dominant effect on performance [bib20]. Bayesian optimization further advances HPO by learning a probabilistic surrogate of the objective function and selecting new configurations via an explicit exploration–exploitation trade-off, enabling faster convergence with significantly fewer evaluations [bib21]. Despite their success, these methods generally treat each optimization problem independently, failing to exploit knowledge from prior tasks  [bib2]. This limitation reduces scalability and sample efficiency, especially in computationally demanding settings such as time-series forecasting, where training costs are high and model behavior is strongly data-dependent [bib3].
@@
To overcome this limitation, meta-learning has emerged as an effective paradigm for improving scalability and resource efficiency across a wide range of learning problems [bib22]. By leveraging knowledge acquired from previously solved tasks, meta-learning enables hyperparameter optimization to adapt more rapidly to new tasks, effectively avoiding repeated optimization from scratch. In particular, Bayesian optimization has been shown to benefit significantly from meta-learning–driven warm starts, where prior task information guides the initial search process  [bib23, bib24].This combination has been successfully applied to complex attention-based architectures,leading to improved forecasting accuracy in energy systems compared to standalone optimization approaches  [bib25].Building on this idea, meta-Bayesian optimization learns informative surrogate model priors from related tasks, further accelerating convergence and improving optimization efficiency on new problems  [bib26].
@@
Beyond Bayesian and meta-learning–based approaches, reinforcement learning (RL) has also been explored for hyperparameter optimization by formulating HPO as a sequential decision-making problem. In this setting, an RL agent iteratively selects hyperparameter configurations based on feedback from previous evaluations, learning a policy that aims to maximize expected performance while reducing the number of costly evaluations  [bib27]. Building on this idea, metalearning has been combined with RL to further improve efficiency. For example, MultiTaskPB2 leverages metadata from prior RL training runs to warm-start hyperparameter optimization and bias surrogate models for new tasks, achieving significantly faster and more effective optimization than standard methods  [bib28].
@@
More recently, large language models (LLMs) have introduced a new paradigm for hyperparameter optimization by leveraging natural language understanding,contextual reasoning, and few-shot generalization capabilities s [bib29, bib30, bib32, bib33]. In AgentHPO  [bib29], an LLM-based agent autonomously analyzes task descriptions and historical trials, proposes new hyperparameter configurations, executes experiments, and iteratively refines its recommendations, forming a human-like and interpretable optimization loop that reduces manual intervention. LLMs have also been integrated with traditional HPO strategies to improve efficiency.For instance, coarse grid search can be guided by an LLM to identify and refine promising regions of the search space, effectively pruning unlikely configurations and reducing computational cost while maintaining or improving tuning quality  [bib32]. Furthermore, LLMs have been combined with Bayesian optimization in frameworks such as LLAMBO, where the BO process is reformulated as a natural language reasoning task. This enables LLMs to propose and evaluate candidate configurations conditioned on past trials, supporting zero-shot warm-starting, improved surrogate modeling, and more focused exploration of promising regions, ultimately leading to faster convergence and superior hyperparameter tuning compared to conventional BO [bib16].Beyond directly steering the optimizationprocess, recent research has investigated how LLMs can be used as meta-learners that reason over collections of prior experiments. Frameworks such as MetaLLMix  [bib17] construct structured meta-databases from historical training runs and rely on LLMs to synthesize transferable optimization knowledge across tasks. Related in-context meta-learning approaches  [bib18] encode datasets and experimental settings as structured, human-readable metadata, enabling LLMs to infer suitable hyperparameter configurations in zero-shot or few-shot regimes.
@@
Unlike agent-based or BO-integrated LLM methods, these approaches emphasize knowledge transfer at the task level rather than iterative trial-by-trial optimization. However, they have primarily been evaluated on classification and regression benchmarks and do not explicitly address the domain-specific challenges of time-series forecasting, where strong temporal dependencies, heterogeneous sampling rates, seasonal dynamics, and high training costs require specialized meta-features and optimization strategies. To address these limitations, we propose LLM-AutoMOpt, a forecasting-specific HPO framework that tightly integrates Bayesian optimization for structured initialization with LLM based in-context reasoning over rich temporal meta-knowledge. In contrast to prior work, LLM-AutoMOpt explicitly accounts for temporal structure and model dynamics, enabling more sample-efficient, interpretable, and adaptive hyperparameter optimization for deep sequence models such as LSTM, Bi-LSTM, and Transformer architectures.
3 Materials and Methods
This section presents the methodology we designed to evaluate the capabilities of LLM-AutoOpt in performing hyperparameter optimization for deep learning models in time-series forecasting. Our goal is to assess how effectively large language models (LLMs) can recommend high-quality hyperparameter configurations while providing interpretable reasoning and leveraging task-specific meta-knowledge. To this end, we first formalize the problem statement in Sect. 3.1. We then describe the construction of forecasting-specific meta-knowledge in Sect. 3.2, followed by the structured prompt design in Sect. 3.4, and finally, the overall LLM-AutoOpt optimization workflow in Sect. 3.5.
3.1 Problem Statement
We consider the problem of hyperparameter optimization (HPO) for deep learning models applied to time-series forecasting. Let $\mathcal{D}=\{(x_{t},y_{t})\}_{t=1}^{T}$ denote a univariate or multivariate time-series dataset with $T$ sequential observations, and let $\mathcal{M}_{\theta}$ represent a forecasting model parameterized by weights $\theta$. The model’s behavior is controlled by a set of hyperparameters $\lambda$ drawn from a search space $\Lambda$, and its predictive performance is evaluated using a loss function $\mathcal{L}(\mathcal{M}_{\theta},\mathcal{D})$.
The HPO task can be formally expressed as the search for the optimal hyperparameter configuration $\lambda^{*}$ that minimizes the expected validation loss:
$\lambda^{*}=\arg\min_{\lambda\in\Lambda}\ \mathbb{E}_{\theta\sim\text{Train}(\mathcal{M}_{\lambda},\mathcal{D})}\Big[\mathcal{L}(\mathcal{M}_{\theta},\mathcal{D}_{\text{val}})\Big],$ (1)
where $\mathcal{M}_{\lambda}$ denotes the model trained under hyperparameters $\lambda$, $\mathcal{D}_{\text{val}}$ is a validation set, and $\text{Train}(\mathcal{M}_{\lambda},\mathcal{D})$ represents the stochastic training procedure that produces model parameters $\theta$.
The primary objective is to efficiently identify $\lambda^{*}$ while providing interpretable insights into the search process. This formulation naturally accounts for the stochastic nature of model training and the high-dimensional, often complex structure of the hyperparameter space, which poses significant challenges in both optimization efficiency and result interpretability.
@@ 3.2 Forecasting-Specific Meta-Knowledge
3.3 Per-Feature Meta-Feature Extraction
@@
3.3.1 Meta-Feature Statistical Summarization
Beyond providing raw meta-features, we derive higher-level statistical summaries that aggregate these descriptors into interpretable indicators of the dataset’s behavior. Examples include summarizing autocorrelation coefficients across multiple lags to characterize overall temporal dependence, aggregating trend and stationarity statistics (e.g., ADF statistics and trend slopes) to describe global non-stationarity, and combining volatility and noise-related measures (e.g., coefficient of variation, residual variance, and outlier ratio) to estimate the effective noise level of the series. These summarized representations are injected into the prompt in human-readable form, enabling the LLM to reason over dataset characteristics at a semantic level rather than directly interpreting raw numerical statistics. This Abstraction facilitates better generalization in decision-making and significantly reduces hallucinated reasoning by constraining the LLM’s inference process to validate dataset-derived indicators.
3.3.2 Bayesian Optimization Initialization
We employ Bayesian Optimization (BO) to efficiently explore the hyperparameter space. Let $\lambda_{\text{BO}}^{\min}$ denote the best configuration found:
$\lambda_{\text{BO}}^{\min}=\arg\min_{\lambda\in\Lambda}\mathcal{L}(\mathcal{M}_{\lambda},\mathcal{D}_{\text{val}})$ (2)
This configuration is provided as meta-knowledge to the LLM, giving it a strong inductive bias.
3.3.3 Model Specification